# Learning High-Degree Parities: The Crucial Role of the Initialization

### v1

@@ E.1 Experiment Details
In the left plot of Figure 2, we are computing the quantity $\mathbb{E}_{w}\left[\mathbb{E}_{x,r}\left[\frac{\partial L(w,x,f(x))}{\partial w_{d}}-\frac{\partial L(w,x,r)}{\partial w_{d}}\right]^{2}\right],$ for $w\sim\mathcal{N}(0,\frac{1}{d}\operatorname{Id}_{d})$, $w\sim\operatorname{Rad}(1/2)$ and $w=\frac{1}{\sqrt{d}}r^{\prime}+\sigma g$ with $r^{\prime}\sim\operatorname{Rad}(1/2)$, and $g\sim\mathcal{N}(0,\frac{1}{d}\operatorname{Id}_{d})$. The function $f$ is the full parity, $r\sim\operatorname{Rad}(1/2)$ and $L(w,x,y):=\max\left(0,1-y\operatorname{ReLU}(w.x)\right)$ is the hinge loss. For the approximated part we update the weights according to $\psi^{t+1}=\psi^{t}-\gamma\left(\Gamma_{r}(\psi^{t})\right)$, with $\psi^{0}\sim\mathcal{N}(0,\frac{1}{d}\operatorname{Id}_{d})$ and $\gamma=1$, and we calculate $\mathbb{E}_{\psi^{t}}\left[\mathbb{E}_{x,r}\left[\frac{\partial L(\psi^{t},x,f(x))}{\partial\psi^{t}_{d}}-\frac{\partial L(\psi^{t},x,r)}{\partial\psi^{t}_{d}}\right]^{2}\right].$
@@ E.2 Additional Experiments
Output Layer Training with Correlation Loss.
The purpose of Figure 5 is to empirically verify our positive theoretical result from Theorem 5. To that purpose, we train a two-layer fully connected network with Rademacher initialization and clipped-ReLU activation $\sigma(x)=\min(1,\max(0,x))$ on the full parity task. We train only the output layer, consistently with our positive result, with SGD with large batch size (1024) with the correlation loss and online fresh samples, until convergence of the test accuracy. We show the test accuracy achieved for different input dimensions ($d$) and different widths of the hidden layer ($w$). Consistently with our theory, with clipped-ReLU $d^{2}$ hidden neurons are sufficient to achieve accuracy $1$.
Figure 5: Learning the full parity with a 2-layer network, where only the output layer is trained by SGD with the correlation loss. We report the test accuracy achieved after training, for clipped-ReLU activation, for different input dimensions ($d$) and hidden layer width ($w$).
@@
Larger input dimension.
In Figure 8, we plot the test accuracy achieved by a 4-layer MLP with varying input dimension $d\in\{100,150,200\}$ and fixed hidden layers widths, trained with the hinge loss on the full parity task, with different $\sigma$-perturbed initializations. We report only the curves for small $\sigma$. We observe that for fixed $\sigma$, learning becomes hard as $d$ increases.

### v3

@@ E.1 Experiment Details
In the left plot of Figure 2, we are computing the quantity $\mathbb{E}_{w}\left[\mathbb{E}_{x,r}\left[\frac{\partial L(w,x,f(x))}{\partial w_{d}}-\frac{\partial L(w,x,r)}{\partial w_{d}}\right]^{2}\right],$ where $w\sim\mathcal{N}(0,\frac{1}{d}\operatorname{Id}_{d})$ for one case and $w\sim\operatorname{Rad}(1/2)$ for the other case, $f$ is the full parity, $r\sim\operatorname{Rad}(1/2)$ and $L(w,x,y):=\max\left(0,1-y\operatorname{ReLU}(w.x)\right)$ is the hinge loss. For the approximated part we update the weights according to $\psi^{t+1}=\psi^{t}-\gamma\left(\Gamma_{r}(\psi^{t})\right)$, with $\psi^{0}\sim\mathcal{N}(0,\frac{1}{d}\operatorname{Id}_{d})$ and $\gamma=1$, and we calculate $\mathbb{E}_{\psi^{t}}\left[\mathbb{E}_{x,r}\left[\frac{\partial L(\psi^{t},x,f(x))}{\partial\psi^{t}_{d}}-\frac{\partial L(\psi^{t},x,r)}{\partial\psi^{t}_{d}}\right]^{2}\right].$
@@ E.2 Additional Experiments
Larger input dimension.
In Figure 5, we plot the test accuracy achieved by a 4-layer MLP trained with the hinge loss on the full parity task, with different $\sigma$-perturbed initializations. We report only the curves for small $\sigma$. We observe that for fixed $\sigma$, learning becomes hard as $d$ increases.
@@
Output Layer Training with Correlation Loss.
The purpose of Figure 9 is to empirically verify our positive theoretical result from Theorem 5. To that purpose, we train a two-layer fully connected network with Rademacher initialization with ReLU and clipped-ReLU activation $\sigma(x)=\min(1,\max(0,x))$ on the full parity task. We train only the output layer, consistently with our positive result, with SGD with large batch size (1024) with the correlation loss and online fresh samples, until convergence of the test accuracy. We show the test accuracy achieved for different input dimensions ($d$) and different widths of the hidden layer ($w$). Consistently with our theory, with clipped-ReLU, $d^{2}$ hidden neurons are sufficient to achieve accuracy $1$ (left). For ReLU, we observe that $w=O(d^{2})$ is not enough to achieve perfect accuracy and we believe that our theoretical bound (i.e $\Omega(d^{4})$) for learning with accuracy $1$ may not be tight and that $d^{3}$ or $d^{3.5}$ may be sufficient (right).
Figure 9: Learning the full parity with a 2-layer network, where only the output layer is trained by SGD with the correlation loss. We report the test accuracy achieved after training, for clipped-ReLU activation (left) and ReLU (right), for different input dimensions ($d$) and hidden layer width ($w$).